OpenStudy (anonymous):

3x^2+8x+4 has what kind of roots?

OpenStudy (anonymous):

two complex?

OpenStudy (anonymous):

or two positive?

OpenStudy (anonymous):

both has to be negative real roots...

OpenStudy (anonymous):

so then it is 2 negative?

OpenStudy (anonymous):

x=-2 is one find the other...

OpenStudy (anonymous):

the other must be positive then right?

OpenStudy (anonymous):

\[b^2-4ac=8^2-4\times 4\times 3=16\]

OpenStudy (anonymous):

has 2 rational roots

OpenStudy (anonymous):

-2 and -2/3 --- both negative